Write the equation in slope-intercept form.
a.)8x-4y=16 b.)x-y=-2 c.)-x+3y=-9

Answer by venugopalramana(3286)

a.)8x-4y=16 slope intercept form of a st.line is Y=mX+c..so rewrite the eqn.as required..
8x-16=4y
4y=8x-16
y=(8/4)x-(16/4)
y=2x-4..which is in the slope intercept form..with slope-m=2 and intercept=c=-4
NOW YOU CAN DO THE OTHER PROBLEMS IN THE SAME WAY.